OpenStudy (anonymous):

a=30cm b=40cm c=? what is c?

OpenStudy (king):

use pythogoras theorem a^2+b^2=c^2

OpenStudy (anonymous):

with some practice you will recognize this as a 3-4-5 right triangle

OpenStudy (anonymous):

with everything multiplied by 10

OpenStudy (anonymous):

c^2= a^2 +b^2 c^2=30^2+40^2 c^2=900+1600 c^2=\[\sqrt{2500}\] c=50 is this correct